# On the Foundation of Distributionally Robust Reinforcement Learning

### v3

@@ 6.2 Stochastic Games
Stochastic games (a.k.a. Markov games) model interactions among players where the environment evolves based on the actions of the players (Solan and Vieille, 2015). In a stochastic game, a group of players participate. During each stage of the game, the game is in a specific state, with each player selecting an action, possibly history-dependent, from a set of possible actions. The actions chosen by the players, along with the current state, dictate the payoff each player receives for that stage and also establish a probability distribution that determines the new state the play will transition to. In our formulation of DRMDP with an S-rectangular adversary, the controller and the adversarial can be viewed as two players with zero-sum rewards, which fits in the framework of stochastic games. This connection is also noted in González-Trejo et al. (2002); Nilim and El Ghaoui (2005); Le Tallec (2007); Shapiro (2021); Li and Shapiro (2023). The line of research concerning stochastic games was pioneered by Shapley (1953), who proved the existence of a stationary equilibrium in a zero-sum game. Following the groundbreaking results of Shapley (1953), a substantial body of literature has emerged, with a primary focus on identifying scenarios in which stationary Markov equilibria exist (Fink, 1964; Takahashi, 1964; Parthasarathy and Sinha, 1989; Nowak, 2003; Simon, 2003; Levy, 2013; Maskin and Tirole, 2001) For a review of the history of stochastic games, we refer readers to Solan and Vieille (2015). Furthermore, stochastic games become a standard framework of multi-agent reinforcement learning (Littman, 1994). We note that the solutions to the DRMDP (2.16) may not be an equilibrium if sup-inf cannot interchange, as there may not exist a consistent belief.
@@
One major difference between stochastic games and DRMDPs lies in the symmetry or asymmetry of the two players. In stochastic games, both players are symmetric in that they can both choose policies that are history-dependent. However, in the DRMDP literature, there is often an asymmetry between the controller and the adversary. Here, the controller’s strategies are history-dependent, while the adversary’s strategies are typically modeled as Markov time-homogeneous. This crucial difference is also noted in Grand-Clement et al. (2023).
DRMDPs with different types of controllers and adversaries are also related to stochastic games with asymmetric information (Nayyar et al., 2013; Nayyar and Gupta, 2017). Their formulations are based on partially observable Markov decision processes and their approach is to reduce the game with asymmetric information to another game with symmetric information.
@@ 7 Conclusion and Future Works
In this paper, we clarify the definitions of distributionally robust Markov decision processes (DRMDP) and identify the scenarios in which the dynamic programming principle (DPP) is either held with full generality or violated. Following this trajectory, several intriguing future directions emerge:
  • In cases where a DPP in the form of the DR Bellman equation does not hold with full generality, are there alternative optimality equations that can characterize the optimal robust control? If such equations exist, can they be effectively employed to facilitate DRRL?
@@ Acknowledgments
The material in this paper is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-20-1-0397. Additional support is gratefully acknowledged from NSF 1915967, 2118199, 2229012, 2312204.
Zhengyuan Zhou graciously acknowledges the generous support provided by NSF 2312205 and NYU Center for Global Economy and Business Grant.
Appendix A Proofs of Theorems in Section 3
In this section, we outline some key properties of of the maxmin values for S-rectangular robust MDPs. We present a proof of Theorem 1 and 2 given these results.
A.1 Proof of Theorem 1
@@
Let $u^{*}$ be the solution to (3.1). Then, $\mu[u^{*}]\geq v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{H}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}),\text{ }\mu[u^{*}]\geq v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{M}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}}),\text{ and }\mu[u^{*}]\geq v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{S}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})$ for all $\mu\in\mathcal{P}(\mathcal{S})$.
@@
Let $u^{*}$ be the solution to (3.1). Then, $\mu[u^{*}]=v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})$ for all $\mu\in\mathcal{P}(\mathcal{S})$.
With these results, we present the proof of Theorem 1 given Lemma 1 and Proposition 3 and 4.
Proof of Theorem 1.

### v4

@@ 7.2 Stochastic Games
Stochastic games (SGs a.k.a. Markov games) model interactions among players where the environment evolves according to their actions [71]. At each stage, the game is at a state, and each player selects an action, potentially history-dependent and randomized, from their decision sets. The realized actions and the current state determine both the players’ payoffs for that stage and the probability distribution transitioning to the next state.
@@
Our formulation of an RMDP with a controller and an S-rectangular adversary can be viewed as a two-player zero-sum game, closely linked to the framework of SGs. This connection has also been observed in González-Trejo et al., [18], Nilim and El Ghaoui, [47], Le Tallec, [34], Shapiro, [59], Li and Shapiro, [37]. Research on SGs was pioneered by Shapley, [64], who proved the existence of a stationary equilibrium in a zero-sum game. Following the groundbreaking results of Shapley, [64], a substantial body of literature has emerged, with a primary focus on identifying scenarios in which stationary Markov equilibria exist [15, 72, 52, 48, 69, 35, 43]. We also note that SGs become a standard framework of multi-agent reinforcement learning [38]. For a review of the history of SGs, we refer readers to Solan and Vieille, [71].
A key distinction between SGs and RMDPs is that solutions to the RMDP Bellman equations (2.14) may not guarantee the existence of a stationary Markov equilibrium when the sup–inf cannot be interchanged. In such cases, the controller can attain a higher value by deviating from the strategy $\pi^{*}$ that achieves the maximum in (3.1) once the adversary’s strategy is fixed as the worst case $\left\{{p_{s}^{*}:s\in S}\right\}$ corresponding to $\pi^{*}$.
@@
Another distinction between RMDPs and lies in the symmetry of the players. In SGs, both players can adopt history-dependent policies, making them symmetric. In contrast, RMDPs often feature asymmetry: the controller may use history-dependent strategies, while the adversary is typically modeled as Markov time-homogeneous, as also noted in Grand-Clement et al., [20].
RMDPs with various controller–adversary attributes are also connected to SGs with asymmetric information [46, 45], which build on partially observable MDPs. These works address the asymmetry by transforming the game into one with symmetric information.
7.3 Distributionally Robust MDPs
Distributionally Robust MDPs (DRMDPs) [80, 10] provide an alternative framework for robust dynamic decision-making. The adversary in a DRMDP can be interpreted from a Bayesian-robust perspective, where it selects a prior distribution over transition probabilities, and the system dynamics are induced by samples drawn from this prior. While DRMDP and RMDP models differ at the level of sample-path dynamics, they are equivalent from the perspective of policy evaluation and optimization. In particular, all of the results in this paper remain valid if we consider the expected version of the adversary. We defer a more detailed discussion to Appendix F.
8 Discussion of Various Attributes
In this section, we provide guidance on selecting the most suitable formulation from the 36 cases presented in Tables 1 to 4 for the DRRL problems. Our criteria for selection are based on four main factors: rectangularity, information structure, convexity of adversaries, and the policy classes of controllers, whether deterministic or randomized. As we shall elaborate in the proceeding discussions, a variety of attributes related to the adversary in the underlying RMDP not only lead to a formulation that facilitates robust decision-making in a dynamic environment but also function as modeling tools to mitigate various forms of model misspecification.
Rectangularity of the adversary:
  • SA-rectangular adversary: SA-rectangular adversary models arise naturally in scenarios where an RL agent is expected to encounter distinct environments or state transitions following different actions. For example, consider an autonomous driving agent deciding whether to turn left, right, or go straight at an intersection. Each of these actions leads to entirely different environments and state observations. In such cases, it is reasonable to assume that the stochasticity associated with state transitions following one action is independent of that associated with other actions. This motivates the use of SA-rectangularity as a natural structural assumption on the adversary in the underlying DRRL formulation. Notably, Ding et al., [14] adopted a DRRL model with a specially structured SA-rectangular ambiguity set and demonstrated that their algorithm outperformed all baselines in autonomous driving tasks.
  • S-rectangular adversary: Compared to SA-rectangular adversaries, S-rectangular adversaries exhibit a more constrained influence by limiting the set of transition probabilities through enforced correlations across different actions. As illustrated in the inventory control Example 1, it is often more natural to restrict the adversary from using the realized action to determine its shift. In effect, S-rectangularity enables the modeler to reduce the adversary’s power by preventing it from reacting to the controller’s chosen action in real time. As a result of this refinement, the resulting robust policies tend to be less conservative. In addition, adding reasonable constraints could facilitate the calibration of the uncertainty set; see, for example, [40].
It is important to note that SA-rectangular adversaries are more powerful than S-rectangular counterparts, which may result in a more conservative policy. Nevertheless, for SA-rectangular adversaries, the existence of a DPP is guaranteed (Table 1), and a $q$-function [78] with its Bellman equation is definable (equation (3.6)). These characteristics facilitate the design of model-based [87, 51, 84, 65, 81, 66, 5] and model-free [39, 76, 77, 83] DRRL algorithms, facilitating efficient learning.
Information structure of the adversary:
  • Markov time-homogeneous adversary: Markov time-homogeneous formulation is widely analyzed in the (R)MDP literature. An adversary with this attribute is compelled to predefine one transition kernel and adhere to its transition probability across all time. Hence, such an adversary is sometimes referred to as static in the literature. Although this formulation is preferred and commonly used by practitioners and DRRL theorists due to its structural simplicity and ease of interpretation [79, 87, 20], the corresponding RMDP problem could pose significant computational challenges due to the absence of a DPP and, consequently, the sub-optimality of time-homogeneous policies. Specifically, DPPs do not exist in general if the ambiguity set is not convex (Table 3) or if the controller is deterministic (Table 4). Therefore, caution is advised for researchers and practitioners when attempting to derive optimal robust control under this formulation. In scenarios where a DPP does not exist, one may develop new exact or approximate methods on a case-by-case basis; alternatively, our asymptotically optimal RL-based policy design (see Section 6) can be consulted for practical guidance in these settings.
    It’s worth noting that previous DRRL papers commonly adopt the modeling perspective of a time-homogeneous adversary [87, 51, 39, 65, 81, 66]. However, these papers all refer to the optimality of deterministic Markov time-homogeneous policies and Bellman equations established in [28], where the principal modeling assumption involves a Markov adversary and a history-dependent controller (as summarized in Table 1). In this paper, we establish the existence or non-existence of the DPP and Bellman equation for both the value and the Q-function in the case of a Markov time-homogeneous adversary, providing the theoretical foundation for these works.
  • Markov adversary: A Markov adversary is the natural modeling choice when the environment is believed to exhibit memoryless transitions under the prescribed state and action spaces. On the other hand, this attribute permits non-stationary transition probabilities. Therefore, a Markov adversary could be an appropriate modeling assumption when the probability structure in the underlying Markov dynamics can vary across time. For instance, in the aforementioned inventory Example 1, the demand distribution could change from day to day, resulting in possible time-nonhomogeneity. From a computational perspective, DPPs also always exist with randomized controllers, see the second columns of Tables 1 and 3.
  • History-dependent adversary: History-dependence allows the adversary to select different distributions based on the evolving history of the system. Even when the modeler has strong reason to believe the environment follows Markov dynamics, partial observability or incomplete modeling of the system state often leads to an optimal controller that is inherently history-dependent, thereby inducing non-Markovian transition dynamics. In this context, formulating the DRRL problem with a history-dependent adversary not only strengthens the robustness of the optimal controller against non-Markov shifts in the environment, but also offers a practical safeguard against model misspecification when the true Markov state is only partially observed or incorrectly specified.
    Moreover, history-dependent adversaries are relevant in RL settings where the true environment is not an MDP, but a DRRL model is adopted for approximation due to computational or data-efficiency considerations. This situation frequently arises when an RL agent interacts strategically with its environment. For instance, consider an RL agent learning bidding strategies in online advertising auctions [6]. In recent years, the industry has shifted from second-price to first-price auction mechanisms [^1] https://blog.google/products/adsense/our-move-to-a-first-price-auction/. In second-price auctions, truthful bidding is a weakly dominant strategy, but this is no longer true under first-price rules. Suppose an RL agent is trained in a second-price environment but is deployed in a first-price exchange. The environment, consisting of strategic competitors, may respond differently to the agent’s varied actions and may incorporate the agent’s bidding history into their strategies. This behavior renders the effective environment non-Markovian.
    In such cases, formulating a fully non-Markovian control problem that explicitly models the strategic learning dynamics of competitors may be prohibitively complex to specify, learn, or solve. In contrast, a DRRL formulation that admits DPP can serve as a computationally tractable and principled approximation.
    Furthermore, from a computational standpoint, a history-dependent adversary model is especially appealing, as the first columns of Tables 14 confirm the guaranteed existence of a DPP.
By symmetry, the controller can also adopt these attributes regarding information availability. The rationale for selecting such information structures for the controller mirrors that of the adversary. Consequently, we do not provide a separate discussion on this point.
Convexity of the adversarial ambiguity sets:
  • Convex adversary: RMDP and DRRL models with convex adversaries are prevalent in literature, with instances spanning various ambiguity sets such as optimal-transport-based [74, 82], convex $f$-divergence-based [84, 66], TV-distance-based [66], as well as the special ambiguity set employed in Xu and Mannor, [80], Wiesemann et al., [79], Ding et al., [14]. Our findings, as presented in Tables 1 and 2, further underscore the computational benefits of convex adversaries. Nevertheless, a word of caution is necessary when navigating scenarios involving a deterministic controller paired with an S-rectangular convex adversary, as the DPP may not be applicable in this case.
  • Non-convex adversary: A non-convex adversary may arise as a modeling choice when the underlying system dynamics are believed to be deterministic [54]. Additionally, the use of a non-convex adversary can result from adopting unions of convex ambiguity sets—often motivated by data-driven calibration procedures that aim to refine the size and geometry of the uncertainty sets, as proposed by Hong et al., [25]. From a computational perspective, the presence of non-convex adversaries not only increases the complexity of solving the robust variants of the Bellman equations (both the value function equation (3.1) and the $q$-function equation (3.4)), but our findings further indicate that the DPP may fail to hold altogether. This breakdown occurs under both deterministic and randomized controller policies, as summarized in Tables 3 and 4.
Possibility of randomization (convexity) of the controller:
  • The randomized controller policy class: While it is well known that deterministic policies can achieve optimality in the non-robust MDP setting (see, e.g. Puterman, [55]), this is generally not the case in the presence of an S-rectangular adversary, a phenomenon first observed and analyzed by Wiesemann et al., [79]. As such, in the context of robust control or DRRL, allowing for randomized controller policies becomes essential for attaining optimal performance. Moreover, from the standpoint of DPP validity, randomized controllers offer an additional advantage: when a history-dependent controller is paired with a Markov adversary, the use of randomized policies can ensure the existence of a DPP (c.f. Tables 3 and 4).
  • The deterministic controller policy class: While optimality may not always be attainable with deterministic policies, implementing randomized policies can be challenging in many real-world scenarios. For example, consider the context of medical treatment planning, where a patient’s current health condition guides the choice among a range of dynamic treatment regimes [7]. In such settings, a randomized policy would require assigning probability distributions over treatment options for a given health state—a practice that is often neither legally permissible nor ethically acceptable.
    When focusing on deterministic policies, it is important to recognize that the absence of a DPP implies that Markov time-homogeneous policies are not guaranteed to be optimal in many cases, even when the adversary’s decision set is convex. This limitation is highlighted in Table 4. To achieve higher rewards in such settings, the controller may need to consider history-dependent policies.
It is worth noting that while our earlier discussion focuses on randomized and deterministic controller policies, the paper extends this distinction to the broader framework of convex versus non-convex controller action sets. Specifically, the randomized controller policy class corresponds to a convex set, while the deterministic policy class is an example of a non-convex set. The DPP results presented in Tables 3 and 4 naturally extend to general convex and non-convex controller policy classes, respectively. From both a modeling and learning perspective, the existence of a DPP for controller policy sets with general geometries opens promising avenues for extending DRRL algorithms. Such extensions could accommodate scenarios in which the controller is restricted to operate within a structured subset of randomized decisions.
@@ 9 Conclusion and Future Works
In this paper, we clarify the definitions of robust Markov decision processes (RMDP) and identify the scenarios in which the dynamic programming principle (DPP) is either held with full generality or violated. Following this trajectory, several intriguing future directions emerge:
  • In cases where a DPP in the form of the robust Bellman equation does not hold with full generality, are there alternative optimality equations that can characterize the optimal robust control? If such equations exist, can they be effectively employed to facilitate DRRL?
@@ Acknowledgments
The material in this paper is partly supported by the Air Force Office of Scientific Research under award number FA9550-20-1-0397. Support from NSF 2229012, 2312204, 2312205, 2403007, 2419564, ONR 13983111, 13983263, and 2025 New York University Center for Global Economy and Business grant is also gratefully acknowledged.
Appendix A Proofs of Auxiliary Results
A.1 Proof of Lemma 1
Proof.
Proof The first inequality follows simply from $\Pi_{\mathrm{H}}\supset\Pi_{\mathrm{M}}\supset\Pi_{\mathrm{S}}$. For the second inequality, recall the inclusion relation (2.10). Thus, for any $\pi\in\Pi=\Pi_{\mathrm{S}},\Pi_{\mathrm{M}},\Pi_{\mathrm{H}}$
This implies the second inequality. ∎
A.2 Proof of Proposition 1
Proof.
ProofWe consider the robust Bellman operator $\mathcal{B}:\mathbb{R}^{S}\rightarrow\mathbb{R}^{S}$ defined as
$\mathcal{B}(u)(s):=\sup_{\phi\in\mathcal{Q}}\inf_{p_{s}\in\mathcal{P}_{s}}E_{\phi,p_{s}}[r(s,A_{0})+\gamma u(X_{1})],\quad\forall s\in S.$
$\left\|\mathcal{B}(u)(s)-\mathcal{B}(v)(s)\right\|_{\infty}$ $\stackrel{{\scriptstyle(i)}}{{\leq}}\max_{s\in S}\sup_{\phi\in\mathcal{Q}}\left|\left(\inf_{p_{s}\in\mathcal{P}_{s}}E_{\phi,p_{s}}[r(s,A_{0})+\gamma u(X_{1})]\right)-\left(\inf_{p_{s}\in\mathcal{P}_{s}}E_{\phi,p_{s}}[r(s,A_{0})+\gamma v(X_{1})]\right)\right|$
$=\gamma\max_{s\in S}\sup_{\phi\in\mathcal{Q}}\left|\inf_{p_{s}\in\mathcal{P}_{s}}E_{\phi,p_{s}}[u(X_{1})]+\sup_{p_{s}\in\mathcal{P}_{s}}-E_{\phi,p_{s}}[v(X_{1})]\right|$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}\gamma\max_{s\in S}\sup_{\phi\in\mathcal{Q}}\max\left\{{\left|\sup_{p_{s}\in\mathcal{P}_{s}}E_{\phi,p_{s}}[u(X_{1})-v(X_{1})]\right|,\left|\inf_{p_{s}\in\mathcal{P}_{s}}E_{\phi,p_{s}}[u(X_{1})-v(X_{1})]\right|}\right\}$
$\leq\gamma\left\|u-v\right\|_{\infty}$
where $(i)$ follows from the inequality $|\sup_{x}f(x)-\sup_{x}g(x)|\leq\sup_{x}|f(x)-g(x)|$ and $(ii)$ is due to $\inf_{x}[f+g]\leq\inf_{x}f+\sup_{x}g\leq\sup_{x}[f+g]$. Therefore, by the Banach fixed point theorem, there exists a unique $u^{*}\in\mathbb{R}^{S}$ s.t. $\mathcal{B}(u^{*})=u^{*}$. Moreover,
A.3 Time-Truncation Technique
To prove the remaining results in the paper, a technique that we will be using in the subsequent proof is to truncate the infinite horizon reward to a finite time $T$ that is large enough. Specifically, consider for any $\pi\in\Pi_{\mathrm{H}}$ and $\kappa\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}$,
$v(\mu,\pi,\kappa)$ $=E_{\mu}^{\pi,\kappa}\left[\sum_{k=0}^{T-1}\gamma^{k}r(X_{k},A_{k})+\gamma^{T}u^{*}(X_{T})\right]+E_{\mu}^{\pi,\kappa}\left[\sum_{k=T}^{\infty}\gamma^{k}r(X_{k},A_{k})-\gamma^{T}u^{*}(X_{T})\right]$ (A.1)
$=E_{\mu}^{\pi,\kappa}\left[\sum_{k=0}^{T-1}\gamma^{k}r(X_{k},A_{k})+\gamma^{T}u^{*}(X_{T})\right]$
$\quad+\frac{\gamma^{T}}{1-\gamma}E_{\mu}^{\pi,\kappa}\left[(1-\gamma)\sum_{k=T}^{\infty}\gamma^{k}r(X_{k},A_{k})-(1-\gamma)\gamma^{T}u^{*}(X_{T})\right]$
$\leq E_{\mu}^{\pi,\kappa}\left[\sum_{k=0}^{T-1}\gamma^{k-T}r(X_{k},A_{k})+\gamma^{k-T}u^{*}(X_{T})\right]+\frac{\epsilon}{2}$
if $\gamma^{T}\leq\epsilon(1-\gamma)/4$. Define
$v_{\mu}^{\pi}(T,\kappa):=E_{\mu}^{\pi,\kappa}\left[\sum_{k=0}^{T-1}\gamma^{k}r(X_{k},A_{k})+\gamma^{T}u^{*}(X_{T})\right],$ (A.2)
where we will use the convention that $\sum_{k=0}^{-1}r_{k}=0$ if $T=0$. In light of (A.1), we will choose $T$ sufficiently large so that $\gamma^{T}\leq\epsilon(1-\gamma)/4$. Then, for all $\pi\in\Pi_{\mathrm{H}}$ and $\kappa\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}$, $|v_{\mu}^{\pi}(T,\kappa)-v(\mu,\pi,\kappa)|\leq\epsilon/2$.
Appendix B Proofs for Dynamic Programming Theorems in Section 3
In this section, we outline some key properties of the maxmin values for S-rectangular robust MDPs. We present a proof of Theorem 1 and 2 given these results.
B.1 Proof of Theorem 1
@@
Let $u^{*}$ be the solution to (3.1). Then, $E_{\mu}[u^{*}(X_{0})]\geq v(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})$, $E_{\mu}[u^{*}(X_{0})]\geq v(\mu,\Pi_{\mathrm{M}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}})$ and $E_{\mu}[u^{*}(X_{0})]\geq v(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})$ for all $\mu\in\mathcal{P}(\mathcal{S})$.
Proposition 2 shows that, without any restrictions on the sets $\mathcal{Q}$ and $\left\{{\mathcal{P}_{s}:s\in S}\right\}$, for any given initial distribution $\mu$, the optimal control values for controllers and adversaries with symmetric information structures are dominated by the $\mu$-expectation of the solution $u^{*}$ of (3.1). On the other hand, by Lemma 1, it is not hard to see that all the optimal control values in Theorem 1 are lower bounded by $v(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})$.
In the next proposition, we show that this smallest value $v(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})$ always equals $E_{\mu}[u^{*}(X_{0})]$.
@@
Let $u^{*}$ be the solution to (3.1). Then, $E_{\mu}[u^{*}(X_{0})]=v(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})$ for all $\mu\in\mathcal{P}(\mathcal{S})$.
With these results, we use Lemma 1 to upper and lower bound all the relevant optimal control values by $E_{\mu}[u^{*}(X_{0})]$ and hence establish Theorem 1.
Proof.